OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept. Show your work.

pooja195 (pooja195):

Hi Welcome to openstudy =) To find the x and y intercept plug in 0 To find x intercept y = -2x - 3 0 = -2x - 3, solve for x To find y intercept plug in 0 for x y = -2x - 3 y = -2(0)- 3